Using the following chemical equation how many grams of water can be produced when 109 L of carbon dioxide are produced?

C4H10 + 5O2 -> 3CO2 + 4H2O


1
Expert's answer
2021-02-15T05:24:26-0500

C4H10 + 5O2 -> 3CO2 + 4H2O

According to the equation, the mass of water equals:

m(H2O) = V(CO2) × 4 × Mr(H2O) / (3 × Vmol) = 109 L × 4 × 18 g/mol / (3 × 22.4 L) = 116.8 g


Answer: 116.8 g